Q.3 Find the Area of the surface generated when the curve x2 + y2 = 1, where y > 0 is revolved about x-axis.

I need this question to be answered ASAP, and I'll give thumbs up and good comments.
Answer #1

(3).Solution: Given curves are x² + y2 = 1, where y20 (1) =y=v1-x?, [:y20] and y=0 If y=0 then from eqn(1): x2 + 02 =1= x=1 y